OpenStudy (anonymous):

The 16-pound turkey cost $14.24. What is the price per pound

OpenStudy (anonymous):

i'm pretty sure it's 89 cents

OpenStudy (texaschic101):

14.24/16 = 0.89 ^^ correct...its 89 cents per lb....because if you multiplied 0.89 * 16, you would get 14.24

OpenStudy (anonymous):

14.24 / 16 = 0.89

OpenStudy (mathstudent55):

Price per pound. Price is in dollars. Pound is the weight, and it is in pounds. "per" suggests division. Price per pounds means the unit cost per pound. Divide the price in dollars by the weight in pounds.

OpenStudy (anonymous):

Thank you everyone

OpenStudy (mathstudent55):

\(\dfrac{$14.24}{16~lb} = 0.89 \dfrac{$}{lb} \) Notice that the units come out correctly in dollars per pound.